OpenStudy (anonymous):

a cylinder has a volume of 1538.6 cubic feet and the height of 10m feet. find the radius

OpenStudy (anonymous):

Pixr^2xh=1538.6 R^2= 1538.6/pix10 R=6.99

OpenStudy (anonymous):

can you give the answer by isolating "???

OpenStudy (anonymous):

i need the answer through isolation

OpenStudy (anonymous):

Pi= 3.142 , V= 1538.6, H=10, R=?

OpenStudy (anonymous):

V= pi x r^2 x h 1538.6= 3.142 x r^2 x 10

OpenStudy (anonymous):

U just have to solve for r

OpenStudy (anonymous):

i know but i need to have a solution by means of isolating. i already know the ans in other solution but i don't know the sol through isolating